# A Paradigm for Creative Ownership

### v1

@@
A Paradigm for Creative Ownership
Tejaswi Polimetla  and  Katy Ilonka Gero Harvard UniversityCambridgeMassachusettsUSA tpolimetla@seas.harvard.edu
(April, 2025; May 2025; May 2025)
@@ A Paradigm for Creative Ownership
copyright: noneccs: Human-centered computing HCI theory, concepts and modelsccs: Social and professional topics Intellectual propertyccs: Human-centered computing Collaborative and social computing theory, concepts and paradigms
*Presented at the Navigating Generative AI Disclosure, Ownership, and Accountability in Co-Creative Domains Workshop at CHIWORK 2025.
@@ 1. Introduction
With the rise of generative AI tools, the determination of ownership in collaborative human-AI settings is a prominent challenge. While legal experts have approached ownership from intellectual property and copyright law, literature in the social sciences and humanities reveals how humans come to form feelings of ownership in the first place. At the same time, while these fields have many conceptualizations of ownership, there does not exist a comprehensive system to diagnose where feelings of ownership come from, why these sentiments arise, and how they can be influenced. In their 2003 paper (Pierce et al., 2003), Pierce et al. define psychological ownership as ”that state where an individual feels as though the target of ownership or a piece of that target is ‘theirs’.” In this paper, we will focus on a narrower definition revolving around creative ownership in which the target of ownership is a creative product or artifact that the individual in question had a role in creating - no matter how small or large. By understanding how humans form feelings of ownership in creative contexts, we hope to form better intuitions in a variety of ownership contexts, as well as find ways to support system designers in preserving human ownership and authenticity. In this paper, we are interested in achieving the following goals: 1) identifying the key criteria required for one to develop a sense of creative ownership over one’s work, and 2) developing an explanatory framework for understanding why different circumstances of creation might lead to different levels of ownership sentiment.
@@ 2.1. Philosophy
One of the most fundamental materialist theories is Locke’s labor theory, which posits that “every man has a property in his own person,” and thereby goes on to argue that when one mixes their labor with natural resources, the resulting good becomes their property - evoking the embodiment theory of personhood (Locke, 1689; Radin, 1982). “Bundle of Rights” views hold ownership as a set of contractual obligations between people in relation to property (Waldron, 1985; Unknown, [n. d.]; Honoré, 1961). More broadly, these theories lend themselves as a basis for the construction of property systems as they tend to be grounded in tangible, enforceable, socially structured claims. In contrast, idealist theories tend to be much more abstract. For instance, Hegel’s ideas of ownership stem from the notion that the ”will” can be embodied in external entities, and that this embodiment is necessary for one’s actualization as a person cannot come to exist without both relation to and differentiation from the external environment (Radin, 1982). While the specifics of theories vary, the investment of one’s self, values, and identity as a means of developing feelings of ownership is a common theme that arises (Sartre, 1943; Émile Durkheim, 1957).
@@ 2.2. Psychology
In addition to these framework based approaches, there has also been much empirical work to understand the emergence of psychological ownership under particular circumstances. In a study of ePortfolio creation in online learning environments, researchers found that psychological ownership is correlated with perceived control of content, planning, personal data and access rights (Buchem, 2012). Levene et al. examined the effects of creation and intent on ownership judged and found that the effects of creation hold even when controlling for other factors. They also showed that successful and intentional creations are ascribed more ownership than unsuccessful or unintentional creations, and that creation is ascribed more ownership than the equivalent labor (Levene and Friedman, 2015). A study of locus of control and psychological ownership found internal locus of control to be positively related to both effectance and self-identity motives (McIntyre et al., 2009).
2.3. Other Areas
In addition to philosophy and psychology, there are a plethora of fields that discuss the emergence of ownership sentiments, specifically in creative contexts. Linguistics literature outlines how people speak about possession and ownership, highlighting how kinship and whole-part relations are indicative of closer links compared to possessor-possessee relations (Aikhenvald, 2012). In her experience as a community dance practitioner, Rosemary Lee details how she focuses on releasing the potential of dancers, with an attention to subtle creative choices in her process (Lee, 2008). In a study of psychological ownership in creative work, Rouse identifies differences in task type (managerial, conceptual, production), task target (abstract, concrete), and task coordination (independent, autonomous) (Rouse, 2013). When measuring ownership sentiments in academic writing contexts, Nicholes found that ownership sentiments tend to be significantly higher in creative genres as opposed to academic genres, and higher for argumentative papers as compared to research papers (Nicholes, 2015). Struggle is another theme that appears, with productive and creative struggle arising as both educational and design goals (Fox, 2023; Zhou and Sterman, 2023).
@@ 3. Framework
Person Process System
@@
Embodiment - How much do you feel that the finished product embodies your values, personality, and identity? Control - How much control did you have over the creative decisions that shaped the final product? Production - How closely did you physically work with final product?
@@
Occupancy - How often do you feel preoccupied with thinking about the product? Intentionality - How intentional were you about the creative decisions that you made? Abstraction - How closely did you conceptually work with the final product?
Recognition - Do you receive a lot of external recognition for your work? Effort - How much physical and/or mental effort did you put into this work? Interdependence - How independent were you during the creative process?
@@
The proposed framework incorporates the material and immaterial components of ownership, tying together elements of individual values and personhood, the creative process, and the systems in which these individuals and processes are embedded. Personhood in a creative context is conceptualized as how the self relates to the creative outcome. How does the creative product embody one’s identity, self and values, how consistently does it occupy the mind of the creator, and to what extent do others recognize the creator (and their personhood) for their work? The creative process is the series of steps by which the creator’s personhood is realized in the creation, and is characterized by the intentionality of the creator, the amount of control they have over the process, and the amount of effort - both mental and physical - that goes into the realization. The system refers to the environment in which this creative process is enacted and its dimensions are borrowed from Rouse’s findings (Rouse, 2013). Abstraction refers to the creator’s role in conceptualizing the work, while production relates to their responsibility for its materialization. Interdependence reflects the creator’s reliance on others or tools to complete the creation. While process and system dimensions may be correlated at times, this is not always the case as shown in Appendix A.
@@
We choose to visualize this framework using three concentric circles, each representing one of the elements listed above, and each element being composed of sub-dimensions. An increase in one dimension leads to an increase not only in the size of its corresponding circle, but in the size of the totality of the three circles altogether. Additionally, this visualization refrains from centering on any one element. Instead, precedence is largely context dependent, and the center-most circle can be thought of as the limiting factor in developing feelings of ownership. There is a threshold slider which is meant to indicate how different individuals may have different thresholds for developing feelings of ownership. The visualization can be explored using this interactive webtool: Creative Ownership Test.

### v2

@@
Abstract.
As generative AI tools become embedded in creative practice, questions of ownership in co-creative contexts are pressing. Yet studies of human-AI collaboration often invoke ”ownership” without definition: sometimes conflating it with other concepts, and other times leaving interpretation to participants. This inconsistency makes findings difficult to compare across or even within studies. We introduce a framework of creative ownership comprising three dimensions - Person, Process, and System - each with three subdimensions, offering a shared language for both system design and HCI research. In semi-structured interviews with 21 creative professionals, we found that participants’ initial references to ownership (e.g., embodiment, control, concept) were fully encompassed by the framework, demonstrating its coverage. Once introduced, however, they also articulated and prioritized the remaining subdimensions, underscoring how the framework expands reflection and enables richer insights. Our contributions include 1) the framework, 2) a web-based visualization tool, and 3) empirical findings on its utility.
@@ Abstract.
copyright: noneccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing HCI theory, concepts and modelsccs: Human-centered computing HCI design and evaluation methods
@@ 1. Introduction
Generative AI tools are now routine parts of a creative practice for many people—from drafting text and composing images to arranging sound and editing video. As teams, platforms, and models co-produce artifacts, creators regularly ask: what does it mean for this to feel ”mine”? While questions of ownership have been asked since antiquity, and collaborations of all kinds can challenge our notions of ownership, generative AI requires us to revisit these questions because these new kinds of ”collaboration” do not map cleanly onto either a human partner nor a simple tool. While law offers formal answers about rights and attribution, legal rights and psychological feelings do not go hand in hand, and Human-Computer Interaction research must grapple with creative ownership as a lived, situated phenomenon.
In their 2003 paper, Pierce et al. (2003) define psychological ownership as ”that state where an individual feels as though the target of ownership or a piece of that target is ‘theirs’.” In this paper, we will focus on a narrower definition revolving around creative ownership in which the target of ownership is a creative product or artifact that the individual in question had a role in creating — no matter how small or large. While prior HCI work engages in challenges involving creative ownership, the concept is not defined consistently. Some studies conflate ownership with adjacent ideas (e.g., control, vision, identity); others elicit participants’ views without a common scaffold, making results hard to compare across settings and media. Existing instruments either center creativity without ownership (e.g., Creativity Support Index (Cherry and Latulipe, 2014)) or synthesize ownership broadly but without a lens for creative work (Kuzminykh and Cauchard, 2020). As a result, researchers and designers lack a compact, comparable way to measure, reason about, and design for creative ownership in human–AI collaboration.
Building upon literature across psychology, philosophy, the humanities and social sciences more broadly, and within human-computer interaction, we introduce a nine-subdimension framework of creative ownership organized across Person, Process, and System. Person captures how the artifact relates to the self; Process characterizes the decisions, intentionality, and effort by which it is created; System situates creation within its material, collaborative, and contextual conditions. We operationalize the framework as an interactive web tool to make the dimensions legible to users, and to standardize comparison across projects and fields.
We conducted semi-structured interviews with 21 creative professionals across a diverse range of fields. We used a two-phase, within-participant protocol. Participants first described one high-ownership and one low-ownership project without the framework, then used our instrument to rate both works and reflect on the output. Qualitatively, pre-framework talk concentrated on a limited subset of subdimensions (embodiment, control, abstraction). Once introduced, participants articulated and prioritized all nine subdimensions, enabling finer distinctions (e.g., conceptual authorship vs. physical production) and revealing medium-dependent nuances. Quantitatively, all nine subdimensions were consistently higher for high-ownership projects than for low-ownership ones; low-ownership ratings showed greater variance, indicating multiple pathways by which ownership can be diminished. Participants reported that the framework was a useful tool for reflection.
This paper makes the following contributions:
  • A concise framework for creative ownership with nine subdimensions that span Person, Process, and System, grounded in psychology, philosophy, and HCI literature.
  • An interactive webtool that enables elicitation, comparison, and visualization of ownership across projects and creative fields.
  • Empirical validation via interviews with 21 creative professionals: (i) quantitative comparison across high and low ownership conditions and (ii) an understanding of how ownership is conceptualized both with and without the presence of the framework.
2. Ownership in the HCI Literature
While considerable work has been done at the intersection of AI and ownership in the HCI literature, there remains little consensus on what ownership entails or how it should be addressed across studies. Researchers often either leave it to participants to define ownership in their own terms, or conflate ownership with adjacent concepts or with a single strand of ownership.
Examples illustrate the breadth of interpretations. One study of artists’ perspectives on AI-generated art focuses primarily on the legal dimensions of ownership, raising questions of disclosure, threats to human labor, and intellectual property (Lovato et al., 2024). In workplace communication, Kadoma et al. (2024) conceptualize ownership as ”mineness”. Draxler et al. (2024) and Liang and Gero (2025) adopt a similar approach, asking questions such as ”to whom should this text belong” or ”I feel like this writing was purely produced by me” in the context of AI-supported writing. Others examine content ownership specifically in the setting of LLM-based writing assistants (Wasi et al., 2024; He et al., 2024). Some researchers employ more structured methods: Joshi and Vogel (2025), for instance, ground their work in psychological ownership, asking about personal ownership, responsibility, personal connection, and emotional attachment. While this captures psychological dimensions, it does not address creativity. Conversely, Wu et al. (2023) adapt the Creativity Support Index, which highlights creative aspects but not ownership.
Other work conflates ownership with related constructs, or dimensions that might be considered only part of ownership. Weber et al. (2025), for example, use the term ”artistic ownership” in studying support for creative goals, yet operationalize it through adjacent concepts such as creative vision, intentions, collaboration, pride, control, and emotional response (Weber et al., 2025). Even when researchers begin with a focused definition, as in Wasi et al.’s work (Wasi et al., 2024) on content ownership, related ideas often surface—embodiment, identity, originality, and effort among them.
Still others leave ownership undefined, inviting participants to articulate their own interpretations. In study of AI-driven scriptwriting by Weber et al. (2024), participants associated ownership with ease, expression, collaboration, uniqueness, and enjoyment. Guo et al.’s work (Guo et al., 2024) on AI value alignment in collaborative ideation found participants emphasizing concept, originality, and creative agency. Similarly, Biermann et al.’s study (Biermann et al., 2022) of AI-supported story writing reported themes such as plot, style, control, concept, and collaboration.
Efforts have been made in HCI to establish more unified frameworks, though these remain limited in scope. Two key examples are Cherry et al.’s Creativity Support Index and Kuzminykh et al.’s synthesis of ownership research in HCI (Cherry and Latulipe, 2014; Kuzminykh and Cauchard, 2020). Cherry and Latulipe (2014) identify six dimensions of creativity support, yet the framework does not establish a direct connection to ownership. Conversely, Kuzminykh and Cauchard (2020) provide a broad overview of ownership literature, but their synthesis is not readily suited to creative contexts. Building on these efforts, our aim is to develop a framework for ownership that is specifically tailored to creative practice and designed for use in HCI research.
@@ 3.1. Philosophy
One of the most fundamental materialist theories is Locke’s labor theory, which posits that ”every man has a property in his own person,” and thereby goes on to argue that when one mixes their labor with natural resources, the resulting good becomes their property - evoking the embodiment theory of personhood (Locke, 1689; Radin, 1982). ”Bundle of Rights” views hold ownership as a set of contractual obligations between people in relation to property (Waldron, 1985; Unknown, [n. d.]; Honoré, 1961). More broadly, these theories lend themselves as a basis for the construction of property systems as they tend to be grounded in tangible, enforceable, socially structured claims. In contrast, idealist theories tend to be much more abstract. For instance, Hegel’s ideas of ownership stem from the notion that the ”will” can be embodied in external entities, and that this embodiment is necessary for one’s actualization as a person cannot come to exist without both relation to and differentiation from the external environment (Radin, 1982). While the specifics of theories vary, the investment of one’s self, values, and identity as a means of developing feelings of ownership is a common theme that arises (Sartre, 1943; Émile Durkheim, 1957).
@@ 3.2. Psychology
In addition to these framework based approaches, there has also been much empirical work to understand the emergence of psychological ownership under particular circumstances. In a study of psychological ownership in creative work, Rouse identifies differences in task type (managerial, conceptual, production), task target (abstract, concrete), and task coordination (independent, autonomous) (Rouse, 2013). In a study of ePortfolio creation in online learning environments, researchers found that psychological ownership is correlated with perceived control of content, planning, personal data and access rights (Buchem, 2012). Levene and Friedman (2015) examined the effects of creation and intent on ownership judged and found that the effects of creation hold even when controlling for other factors. They also showed that successful and intentional creations are ascribed more ownership than unsuccessful or unintentional creations, and that creation is ascribed more ownership than the equivalent labor. A study of locus of control and psychological ownership found internal locus of control to be positively related to both effectance and self-identity motives (McIntyre et al., 2009). When measuring ownership sentiments in academic writing contexts, Nicholes (2015) found that ownership sentiments tend to be significantly higher in creative genres as opposed to academic genres, and higher for argumentative papers as compared to research papers. Struggle is another theme that appears, with productive and creative struggle arising as both educational and design goals (Fox, 2023; Zhou and Sterman, 2023).
@@ 4. Framework
Our framework consists of nine dimensions of creative ownership:
  • Person:
@@
    • Embodiment – How much do you feel that the finished product embodies your values, personality, and identity?
@@
    • Occupancy – How often do you feel preoccupied with thinking about the product?
    • Recognition – Do you receive a lot of external recognition for your work?
  • Process:
    • Control – How much control did you have over the creative decisions that shaped the final product?
    • Intentionality – How intentional were you about the creative decisions that you made?
    • Effort – How much physical and/or mental effort did you put into this work?
  • System:
    • Production – How closely did you physically work with the final product?
    • Abstraction – How closely did you conceptually work with the final product?
    • Interdependence – How independent were you during the creative process?
@@
The proposed framework incorporates the material and immaterial components of ownership, tying together elements of individual values and personhood, the creative process, and the systems in which these individuals and processes are embedded. Personhood in a creative context is conceptualized as how the self relates to the creative outcome. How does the creative product embody one’s identity, self and values, how consistently does it occupy the mind of the creator, and to what extent do others recognize the creator (and their personhood) for their work? The creative process is the series of steps by which the creator’s personhood is realized in the creation, and is characterized by the intentionality of the creator, the amount of control they have over the process, and the amount of effort - both mental and physical - that goes into the realization. The system refers to the environment in which this creative process is enacted and its dimensions are borrowed from Rouse’s findings (Rouse, 2013). Abstraction refers to the creator’s role in conceptualizing the work, while production relates to their responsibility for its materialization. Interdependence reflects the creator’s reliance on others or tools to complete the creation. While process and system dimensions may be correlated at times, this is not always the case as shown in Appendix A, which illustrates how these dimensions can vary through a series of case studies.
@@
We choose to visualize this framework using three concentric circles, each representing one of the elements listed above, with each element composed of sub-dimensions. An increase in one dimension leads not only to an increase in the size of its corresponding circle, but also in the total size of the three circles together. Importantly, the visualization refrains from centering on any one element. Instead, precedence is context dependent, and the innermost circle can be thought of as the limiting factor in developing feelings of ownership. A threshold slider illustrates how individuals may have different thresholds for developing such feelings. The visualization can be explored using this interactive webtool (Figure 1): Creative Ownership Test.
Figure 1. Creative Ownership Test Webtool